OpenStudy (anonymous):

Determine whether the series convergent or divergent (sqrt(n)+4)/n^2? I'm at a loss for this one. Please do not just tell me "this series converges." I know it does. Is this just a p-series test?

zepdrix (zepdrix):

\[\Large\rm \sum_{n=1}^{\infty}\frac{\sqrt{n}+4}{n^2}\ge\sum_{n=1}^{\infty}\frac{\sqrt{n}}{n^2}\qquad \forall ~n\ge1\]Hmmm comparison test maybe? Seems like this should work. Yes, relating it to P-series as you suggested.

OpenStudy (zarkon):

\[0\le\frac{\sqrt{n}+4}{n^2}\le\frac{\sqrt{n}+4\sqrt{n}}{n^2}\le\frac{5\sqrt{n}}{n^2}\] for \(n\ge 1\)

OpenStudy (zarkon):

I guess I could have put \[\frac{\sqrt{n}+4\sqrt{n}}{n^2}=\frac{5\sqrt{n}}{n^2}\] but it is still technically true

zepdrix (zepdrix):

Oh we want to show that our series is less than a convergent series :3 lol ya i was thinking backwards XD

OpenStudy (anonymous):

I got it! Thank you very much!